49 only if you or less.
I just want th final answerI just want th final answer
I just want th final answer
I just want th final answer
Answer #1

Nowo panticke is unden gravity 2-43 しに。 8.8仁缸ss 85 七= | 2x8.PS S 8 七_smee Vre 6-3Ph t = 6.57 0.335 kc 9-8 2x9.8 3.23213 ん= 3.